SpaceRef

SpaceRef


NOAA SATOPS Morning Report: Friday, November 18, 2011

Status Report From: NOAA Office of Satellite Operations
Posted: Friday, November 18, 2011

Customers or users who want to call to ask about problems or report outages can use the following contacts:

- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 322

GOES
All GOES operations were nominal over the past 24 hours.
No aother nomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
N19 DOY321 06:13z Partial (11242 of 13800 frames) GAC 3A transferred to EUMETSAT due

to a scheduling problem.

All other POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety

ESPC Reported data outage and delays
Due to a ESPC server problem at NSOF, the following GOES West sounder products were

unavailable from 18z to 20z; specifically the 1800Z West LCB/ASH Conus, Alaska and

Hawaii, 1824Z GOES West Imager ASOS, 1800Z Imager DPI West, 1830Z to 2000Z GOES

West ABBA, and the 1745z and 1800Z CSBT.

Questions or comments to Ron.Mahmot@noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.